If the perimeter of a rectangular field is 60 meters. Its area is 200 square meters. Find its dimensions.
Answer by Cromlix(4381)

Using x for length and y for breadth
Perimeter = 2x + 2y = 60
Area = xy = 200
x = 200/y
Substitute x = 200/y
in 2x + 2y = 60
2(200/y)+ 2y = 60
200/y + 2y - 60 = 0
Take to a common denominator
= y
200 + 2y^2 - 60y = 0
rearrange
2y^2 - 60y + 200
2( y^2 - 30y + 100)
2(y - 10)(y - 20)
Substitute y = 10 and y = 20
into 2x + 2y = 60
2x + 2(10) = 60
2x + 20 = 60
2x = 40
x = 20
2x + 2(20) = 60
2x + 40 = 60
2x = 20
x = 10
So either the length is 10 m and breadth is 20m
or
length is 20m and breadth is 10m